# From Language to Cognition: How LLMs Outgrow the Human Language Network

### v1

@@ Methods
We first ask if the alignment procedure is meaningful, i.e., whether the encoding models capture meaningful linguistic information and generalize to new linguistic contexts. Figure 2(a) shows average brain alignment across all brain datasets under three conditions: (1) a pretrained model processing original stimuli, (2) a pretrained model processing random token sequences, and (3) an untrained model processing original stimuli. To evaluate metric reliability, we expect random sequences to yield significantly lower alignment than real stimuli. However, CKA fails this criterion, assigning similar alignment scores to both, and even untrained models surpass pretrained ones. In contrast, linear predictivity correctly differentiates between real and random stimuli, with a stronger separation than RSA.
@@
The second criterion we propose is that LLMs with high brain alignment should be able to generalize to held-out stimuli, with a preference for generalizing far outside the stimuli used for mapping model to brain activity. A key factor in designing a corresponding cross-validation scheme is contextualization—how the data is split into train and test sets Feghhi et al. (2024). The Pereira2018 dataset consists of 24 topics composed of multi-sentence passages, and sentences are presented in their original order to both humans and models. A random sentence split (contextualization) allows sentences from the same topic in both train and test sets, and is thus less demanding of generalization. A stronger generalization test ensures entire topics are held out, preventing models from leveraging shared context. Figure 2(b) shows that contextualization makes it easier for model to predict brain activity. In contrast, topic-based splits halve the raw alignment score for pre-trained models. The score of untrained models is reduced even more strongly when enforcing generalization across topics, suggesting that much of their alignment is context-dependent. Nonetheless, untrained models retain significant alignment – about half of pre-trained models – even with strong generalization requirements. We next explore the contributing factors for the alignment of models with untrained parameters.
Figure 4: Brain Alignment Saturates Early on in Training. Plots indicate the brain alignment scores of six models from the Pythia model suite with varying sizes (log x-axis up to 16B tokens, uneven spacing after black line). Scores are normalized by their cross-subject consistency scores. Alignment quickly peaks around 2–8B tokens before saturating or declining, regardless of model size.
What Makes Untrained Models Brain Aligned
@@
We evaluate the brain alignment of various LLMs with untrained parameters to determine which architecture exhibits the strongest inductive bias toward the human language network. Figure 3(a) presents the average alignment across five different random initializations for each untrained model. Each model consists of a stack of two building blocks from its respective architecture, with a hidden state of $1024$. To ensure a fair comparison, we apply the localizer to the output representations of the last token in the sequence from these two blocks, extracting 128 units to predict brain activity.
@@
Our findings reveal two key insights. First, sequence-based models—such as GRU, LSTM, Transformers, and even a simple mean operation over token representations—exhibit higher brain alignment than models that rely solely on the last token’s representation, such as Linear or MLP. In other words, context or temporal integration is a crucial factor in achieving high alignment. Second, we observe a notable difference between Transformer-v1 and Transformer-v2. While Transformer-v2 applies static positional embeddings by directly adding them to token embeddings, Transformer-v1 uses rotary position encoding. Our results suggest that static positional encoding enables models to capture intrinsic temporal dynamics in sentences, providing further evidence that temporal integration is critical for brain-like language representations.
@@
Figure 5: Formal Competence Tracks Brain Alignment More Closely Than Functional Competence. Each column compares how the evolution of formal competence (top) and functional competence (bottom) tracks the evolution of brain alignment during training. Formal and functional competence correspond to low-level and high-level linguistic processing, respectively. The $R^{2}$ values quantify the strength of this relationship, with higher values in formal competence suggesting it as the key driver of the observed brain alignment. (a): The data averaged across models of five different sizes. (b-d): the same comparison as in (a), but with comparisons were made for models from the Pythia suite with three different sizes.
Isolating Critical Components of Transformers
@@
We perform an ablation study on the architectural components of Transformer-v2 using a single block to further isolate the key elements responsible for brain alignment of untrained parameter models (Figure 3(c)). To isolate the architectural components, we study the untrained model. The architectural components analyzed are labeled on the left of each bar in Figure 3(b). AyAttn refers to all components inside the lower box in Figure 3(c), including the first layer norm, multi-head attention, and the residual connection that follows. AyMLP corresponds to the components in the upper box, comprising the post-attention layer norm, MLP, and the subsequent residual layer. AyPos represents the addition of positional embeddings to token embeddings. AyTokens means the model directly returns the raw token embeddings without further processing. This structured ablation helps pinpoint the specific components that contribute most to brain alignment. Once again, we observe that integration across tokens, via attention mechanisms and positional encoding, yields the highest brain alignment. Further, we found that untrained parameter models perform better than chance-level performance on formal competence benchmarks, mirroring their non-zero brain alignment. In contrast, functional competence benchmarks remain at chance level for untrained models. This further supports the finding that brain alignment is primarily driven by formal, rather than functional, linguistic competence. (Figure 3(d)).
Role of Weight Initialization
Using the Transformer-v2 architecture from Figure 3(a), we varied the standard deviation used to initialize the weights of the untrained model (see Figure 9 in the Appendix). Our results reveal that the initialization value yielding the highest alignment is the default setting used in the HuggingFace Transformers library Wolf et al. (2019) suggesting that brain alignment may serve as a useful heuristic for optimizing model initialization. (see also Hosseini et al. (2024)).
How Does Alignment Change Over Training?
Having established what makes an untrained model brain-aligned in the previous section, we now investigate how brain alignment evolves during training. To do this, we use the Pythia model suite Biderman et al. (2023), which consists of models of various sizes, all trained on the same  300B tokens, with publicly available intermediate checkpoints for analysis.
Brain Alignment Evolution During Training
Figure 4 illustrates the brain alignment of six Pythia models across five brain recording datasets at 34 training checkpoints, spanning approximately 300B tokens. Each panel presents checkpoints that are logarithmically spaced up to the vertical line, emphasizing the early-stage increase in brain alignment, which occurs within the first 5.6% of training time. Beyond this point, the panels display the remaining training period, where brain alignment stabilizes. More specifically, we observe the following trend: (1) Brain alignment is similar to the untrained model until approximately 128M tokens. (2) A sharp increase follows, peaking around 8B tokens. (3) Brain alignment then saturates for the remainder of training. Despite the vast difference in model scales shown in Figure 4, the trajectory of brain alignment is remarkably similar.
Brain Alignment Tracks Formal Competence
Following the observation that brain alignment plateaus early in training, we next investigate how this relates to the development of formal and functional linguistic competence in LLMs. Figure 5 displays the average brain alignment alongside the average performance on formal competence benchmarks (top row) and functional competence benchmarks (bottom row). This is shown for three Pythia models (1B, 2.8B, and 6.9B parameters) and the average of five Pythia models (last column) across the training process. To quantify this relationship, we trained a ridge regression model (with a single scalar weight) to predict brain alignment scores from benchmark scores using 10-fold cross-validation. The average R-squared value across these folds serves as our metric for comparing the relationship between formal/functional linguistic competence and brain alignment. These R-squared values are shown in each panel of Figure 5. Finally, we performed a Wilcoxon signed-rank test on the distributions of R-squared values. This revealed that formal competence is significantly more strongly correlated with brain alignment than functional linguistic competence (W = 0.0, p $<$ 0.002). One possible explanation for why brain alignment emerges before formal linguistic competence is that existing LLM benchmarks assess performance using discrete accuracy thresholds (hard metrics), rather than capturing the gradual progression of competence through more nuanced, continuous measures (soft metrics) (Schaeffer et al., 2023).
We show the individual benchmark scores across all checkpoints in Figure 8.
@@
Figure 6: NWP and behavioral alignment correlate with brain alignment only in early training. (Top Row): Correlation between brain alignment and language modeling loss shows a strong, significant relationship during early training (up to 2B tokens). While this correlation weakens in later stages (up to  300B tokens). Results are shown for three models and the average of all 8 models (last column). (Bottom Row): The same analysis, but for the correlation between brain alignment and behavioral alignment, revealing a similar trend—strong correlation early in training, but no significant relationship as models surpass human proficiency.
Superhuman LLMs Lose Behavioral Alignment
@@
Human language processing is strongly modulated by prediction, with words that violate predictions leading to longer reading times (Smith and Levy, 2013). Early in training, LLMs align with this pattern, but as they surpass human proficiency (Shlegeris et al., 2022), their perplexity drops and they begin encoding statistical regularities that diverge from human intuition (Oh and Schuler, 2023; Steuer et al., 2023). This shift correlates with a decline in behavioral alignment, suggesting that superhuman models rely on different mechanisms than those underlying human language comprehension. Figure 6 shows that brain alignment initially correlates with perplexity and behavioral alignment, but only during the early stages of training (up to  2B tokens). Beyond this point, as models surpass human language proficiency, these correlations diminish. In larger models, we observe a negative correlation between brain alignment and behavioral alignment in the later stages of training. This trend reinforces the idea that early training aligns LLMs with human-like processing as also observed in earlier, while in later stages their language mechanisms diverge from humans.
Model Size Does Not Predict Alignment
Figure 7: Model Size Does Not Predict Brain Alignment when localizing a fixed set of language units. Brain alignment across model sizes in the Pythia suite, measured at their final training checkpoints. Brain alignment is shown for each dataset, along with the average alignment across datasets, for eight models of varying sizes.
Figure 7 presents the brain alignment for each dataset, along with the average alignment across datasets, for eight models of varying sizes from the Pythia model suite (final checkpoint). Contrary to the assumption that larger models exhibit higher brain alignment Aw et al. (2023), we observe a decline in average alignment starting from 1B parameters up to 6.9B parameters, when controlling for feature size. This analysis is made possible by functional localization, which allows us to extract a fixed number of units from each model, rather than relying on hidden state dimensions, as done in previous studies. This approach ensures a fairer comparison among models. We show in the Appendix that increasing the number of localized units has minimal impact on the relative ranking of the models. Additionally, these findings align with expectations in the neuroscience language community, where it is widely believed that human language processing does not require superhuman-scale language models to capture neural activity in the brain’s language network.
Conclusion & Future Work
@@
In this work, we investigated how brain alignment in LLMs evolves throughout training, revealing different learning processes at play. We demonstrated that alignment with the human language network (LN) primarily correlates with formal linguistic competence Mahowald et al. (2024), peaking and saturating early in training. In contrast, functional linguistic competence, which involves world knowledge and reasoning, continues to grow beyond this stage. These findings suggest that the LN primarily encodes syntactic and compositional structure, in line with the literature of language neuroscience Fedorenko et al. (2024), while broader linguistic functions may rely on other cognitive systems beyond the LN. We also showed that model size is not a reliable predictor of brain alignment when controlling for feature size. Instead, alignment is determined by architectural inductive biases, token integration mechanisms, and training dynamics. Our standardized brain-scoring framework eliminates contextualization biases from previous work, ensuring more rigorous evaluations. Finally, we demonstrated that current brain alignment benchmarks are not saturated, indicating that LLMs can still be improved in modeling human language processing. Together, these findings challenge prior assumptions about how alignment emerges in LLMs and provide new insights into the relationship between artificial and biological language processing.
Future Work
@@
While this study offers a comprehensive analysis of brain alignment in LLMs, several open questions remain. If functional competence extends beyond the language network, future work should explore which additional brain regions LLMs align with as they develop reasoning and world knowledge, particularly in other cognitive networks like the multiple demand (Duncan and Owen, 2000) or theory of mind network (Saxe and Kanwisher, 2003; Saxe and Powell, 2006). Our findings suggest that LLM brain alignment studied should be broadened from the LN to downstream representations underlying other parts of cognition. This raises the question of whether specific transformer units specialize in formal vs. functional linguistic competence (AlKhamissi et al., 2024; Kumar et al., 2024). A key question remains: Does LLM alignment evolution mirror human language acquisition? Comparing it to developmental data could reveal insights into their learning trajectories, and help differentiate formal from functional language learning. Expanding brain-scoring benchmarks and incorporating multimodal models will help address these questions, further bridging the gap between artificial and biological intelligence and deepening our understanding of how both systems process and represent language.

### v2

@@ 4.1 Robust Metrics and Generalization Tests
We first ask if the alignment procedure is meaningful, i.e., whether the encoding models capture meaningful linguistic information and generalize to new linguistic contexts. Figure 6(a) in Appendix B shows average brain alignment across all brain datasets under three conditions: (1) a pretrained model processing original stimuli, (2) a pretrained model processing random token sequences, and (3) an untrained model processing original stimuli. To evaluate metric reliability, we expect random sequences to yield significantly lower alignment than real stimuli. However, CKA fails this criterion, assigning similar alignment scores to both, and even untrained models surpass pretrained ones. In contrast, linear predictivity differentiates between real and random stimuli, more so than RSA.
@@
The second criterion we propose is that LLMs with high brain alignment should be able to generalize to held-out stimuli, with a preference for generalizing far outside the stimuli used for mapping the model to brain activity. A key factor in designing a corresponding cross-validation scheme is contextualization—how the data is split into train and test sets Feghhi et al. (2024). The Pereira2018 dataset consists of 24 topics composed of multi-sentence passages, and sentences are presented in their original order to both humans and models. A random sentence split (contextualization) allows sentences from the same topic in both train and test sets, and is thus less demanding of generalization. A stronger generalization test ensures entire topics are held out, preventing models from leveraging shared context. Figure 6(b) shows that contextualization makes it easier for the model to predict brain activity. In contrast, topic-based splits halve the raw alignment score for pretrained models. The score of untrained models is reduced even more strongly when enforcing generalization across topics, suggesting that much of their alignment is context-dependent. Nonetheless, untrained models retain significant alignment – about 50% of pretrained models – even with strong generalization requirements.
Figure 3: Brain Alignment Saturates Early on in Training. Plots indicate the brain alignment scores of three models from the Pythia model suite with varying sizes (log x-axis up to 16B tokens, uneven spacing after black line). Scores are normalized by their cross-subject consistency scores. Alignment quickly peaks around 2–8B tokens before saturating or declining, regardless of model size (see Appendix D and F for more models).
5 Results
The following sections progressively unpack the emergence and limits of brain alignment with the human language network in LLMs. Section 5.1 establishes the foundation by showing that untrained models already exhibit modest brain alignment, pointing to the role of architectural priors. Building on this, Section 5.2 tracks how alignment evolves with training and reveals that it strongly correlates with the early acquisition of formal linguistic competence, but less so with functional abilities. Section 5.3 then shows that as models exceed human-level performance in next-word prediction, their brain and behavioral alignment begins to diverge, suggesting that at this point, LLMs outgrow their initial alignment with human language processing.
5.1 Brain Alignment of Untrained Models
@@
We evaluate the brain alignment of various LLMs with untrained parameters to determine which architecture exhibits the strongest inductive bias toward the human language network. Figure 2(a) presents the average alignment across five different random initializations for six different untrained models. Each model consists of a stack of two building blocks from its respective architecture, with a hidden state of $1024$. To ensure a fair comparison, we apply the localizer to the output representations of the last token in the sequence from these two blocks, extracting 128 units to predict brain activity. Our findings reveal two key insights. First, sequence-based models—such as GRU, LSTM, Transformers, and even a simple mean operation over token representations—exhibit higher brain alignment than models that rely solely on the last token’s representation, such as Linear or MLP. In other words, context or temporal integration is a crucial factor in achieving high alignment. Second, we observe a notable difference between Transformer-v1 and Transformer-v2. While Transformer-v2 applies static positional embeddings by directly adding them to token embeddings, Transformer-v1 uses rotary position encoding. Our results suggest that static positional encoding enables models to capture intrinsic temporal dynamics in sentences—possibly tracking evolving word positions—providing further evidence that temporal integration is critical for brain-like language representations.
@@
Figure 4: Formal Competence Tracks Brain Alignment More Closely Than Functional Competence. Each column compares how the evolution of formal competence (top) and functional competence (bottom) tracks the evolution of brain alignment during training. The $R^{2}$ values quantify the strength of this relationship, with higher values in formal competence suggesting it as the key driver of the observed brain alignment. (a): The data averaged across models of five different sizes. (b-d): the same comparison as in (a), but with comparisons were made for models from the Pythia suite with three different sizes.
Key Components of Transformers
@@
To further isolate the key elements responsible for brain alignment in untrained parameter models, we perform an ablation study on the architectural components of Transformer-v2 using a single block (Figure 2(c)). By focusing on the untrained model, we isolate the effect of architecture alone, without confounding influences from training. The architectural components analyzed are labeled on the left of each bar in Figure 2(b). AyAttn refers to all components inside the lower box in Figure 2(c), including the first layer norm, multi-head attention, and the residual connection that follows. AyMLP corresponds to the components in the upper box, comprising the post-attention layer norm, MLP, and the subsequent residual layer. AyPos represents the addition of positional embeddings to token embeddings. AyTokens means the model directly returns the raw token embeddings without further processing. This systematic ablation helps pinpoint the components that contribute most to brain alignment. Once again, we observe that integration across tokens, via attention mechanisms and positional encoding, yields the highest brain alignment. Further, we found that untrained parameter models perform better than chance-level performance on formal competence benchmarks, mirroring their non-zero brain alignment. In contrast, functional competence benchmarks remain at chance level for untrained models. This further supports the finding that brain alignment is primarily driven by formal, rather than functional, linguistic competence. (see Figure 2(d)).
@@
Figure 5: NWP and Behavioral Alignment Correlate with Brain Alignment Only in Early Training. (Top Row): Correlation between brain alignment and language modeling loss shows a strong, significant relationship during early training (up to 2B tokens). While this correlation weakens in later stages (up to ~300B tokens). Results are shown for three models and the average of all 8 models (last column). (Bottom Row): The same analysis, but for the correlation between brain alignment and behavioral alignment, revealing a similar trend—strong correlation early in training, but no significant relationship as models surpass human proficiency.
5.2 Brain Alignment Over Training
Having established the architectural components that make an untrained model brain-aligned in the previous section, we now investigate how brain alignment evolves during training. To do so, we use the Pythia model suite Biderman et al. (2023), which consists of models of various sizes, all trained on the same $\sim$300B tokens, with publicly available intermediate checkpoints. We report results for a model from a different family, SmolLM2-360M (Allal et al., 2025), which provides checkpoints at 250B-token intervals, in Appendix F.
Figure 3 illustrates the brain alignment of six Pythia models across five brain recording datasets at 34 training checkpoints, spanning approximately 300B tokens. Each panel presents checkpoints that are logarithmically spaced up to the vertical line, emphasizing the early-stage increase in brain alignment, which occurs within the first 5.6% of training time. Beyond this point, the panels display the remaining training period, where brain alignment stabilizes. More specifically, we observe the following trend: (1) Brain alignment is similar to the untrained model until approximately 128M tokens. (2) A sharp increase follows, peaking around 8B tokens. (3) Brain alignment then saturates for the remainder of training. Despite the vast difference in model sizes shown in Figure 3, the trajectory of brain alignment is remarkably similar.
Alignment Tracks Formal Competence
Following the observation that brain alignment plateaus early in training, we next investigate how this relates to the emergence of formal and functional linguistic competence in LLMs. Figure 4 displays the average brain alignment alongside the average performance on formal competence benchmarks (top row) and functional competence benchmarks (bottom row). This is shown for three Pythia models (1B, 2.8B, and 6.9B parameters) and the average of five Pythia models (first column) across the training process. To quantify this relationship, we train a ridge regression model (with a single scalar weight) to predict brain alignment scores from benchmark scores using 10-fold cross-validation. The average R-squared value across these folds serves as our metric for comparing the relationship between formal/functional linguistic competence and brain alignment. These R-squared values are shown in each panel of Figure 4. Finally, we perform a Wilcoxon signed-rank test on the distributions of R-squared values. This test reveals that formal linguistic competence is significantly more strongly correlated with brain alignment than functional competence (W = 0.0, p << 0.002). One possible explanation for why brain alignment emerges before formal linguistic competence is that existing LLM benchmarks assess performance using discrete accuracy thresholds (hard metrics), rather than capturing the gradual progression of competence through more nuanced, continuous measures (soft metrics) (Schaeffer et al., 2023). We show the individual benchmark scores across all checkpoints in Figure 8 in Appendix E.
5.3 LLMs Lose Behavioral Alignment
Do language models that improve in next-word prediction remain aligned with human behavioral and neural responses, or do they diverge as they surpass human proficiency? To answer this question we use the Futrell2018 benchmark, which has been widely used in previous research to measure linguistic behavior (Futrell et al., 2018; Schrimpf et al., 2021; Aw et al., 2023). This dataset consists of self-paced reading times for naturalistic story materials from 180 participants. Per-word reading times provide a measure of incremental comprehension difficulty, a cornerstone of psycholinguistic research for testing theories of sentence comprehension (Gibson, 1998; Smith and Levy, 2013; Brothers and Kuperberg, 2021; Shain et al., 2024). We measure alignment by calculating the Pearson correlation between a model’s cross-entropy loss for a specific token in the sequence and the average human per-word reading time. The loss for words that comprise multiple tokens is added together before computing the correlation.
@@ 5.3 LLMs Lose Behavioral Alignment
Early in training, LLMs align with this pattern, but as they surpass human proficiency (Shlegeris et al., 2022), their perplexity drops and they begin encoding statistical regularities that diverge from human intuition (Oh and Schuler, 2023; Steuer et al., 2023). This shift correlates with a decline in behavioral alignment, suggesting that superhuman models rely on different mechanisms than those underlying human language comprehension. Figure 5 shows that brain alignment initially correlates with perplexity and behavioral alignment, but only during the early stages of training (up to ~2B tokens). Beyond this point, these correlations diminish. In larger models, we observe a negative correlation between brain alignment and behavioral alignment in the later stages of training. This trend reinforces that early training aligns LLMs with human-like processing as also observed in earlier stages, while in later stages their language mechanisms diverge from humans.
6 Conclusion
In this work, we investigate how brain alignment in LLMs evolves throughout training, revealing different learning processes at play. We demonstrate that alignment with the human language network (LN) primarily correlates with formal linguistic competence Mahowald et al. (2024), peaking and saturating early in training. In contrast, functional linguistic competence, which involves world knowledge and reasoning, continues to grow beyond this stage. These findings suggest that the LN primarily encodes syntactic and compositional structure, in line with the literature of language neuroscience Fedorenko et al. (2024a), while broader linguistic functions may rely on other cognitive systems beyond the LN. This developmental approach reveals when brain-like representations emerge, offering a dynamic perspective compared to prior work focused on fully trained models. For example, Oota et al. (2023) demonstrated that syntactic structure contributes to alignment by selectively removing specific properties from already trained models. In contrast, we show that formal linguistic competence actively drives brain alignment during the early phases of training. Similarly, Hosseini et al. (2024) reported that models achieve strong alignment with limited data; we identify why: the brain-like representations emerge as soon as core formal linguistic knowledge is acquired. Further, their study evaluated only four training checkpoints and 2 models on a single dataset (Pereira2018). Our study evaluated eight models (14M–6.7B parameters) across 34 checkpoints spanning 300B tokens, and used five neural benchmarks within a rigorous brain‑scoring framework. This extensive design enabled fine‑grained correlations with both formal and functional linguistic benchmarks and ensured our results are robust and generalizable.
@@ 6 Conclusion
We also show that model size is not a reliable predictor of brain alignment when controlling for the number of features (see Appendix I). Instead, alignment is shaped by architectural inductive biases, token integration mechanisms, and training dynamics. Our standardized brain-scoring framework eliminates contextualization biases from previous work, ensuring more rigorous evaluations. Finally, we demonstrate that current brain alignment benchmarks are not saturated, indicating that LLMs can still be improved in modeling human language processing. Together, these findings challenge prior assumptions about how alignment emerges in LLMs and provide new insights into the relationship between artificial and biological language processing.
Limitations
@@
While this study offers a comprehensive analysis of brain alignment in LLMs, several open questions remain. If functional competence extends beyond the language network, future work should explore which additional brain regions LLMs align with as they develop reasoning and world knowledge, particularly in other cognitive networks like the multiple demand (Duncan and Owen, 2000) or theory of mind network (Saxe and Kanwisher, 2003; Saxe and Powell, 2006). Our findings suggest that LLM brain alignment studies should be broadened from the LN to downstream representations underlying other parts of cognition. This raises the question of whether specific transformer units specialize in formal vs. functional linguistic competence (AlKhamissi et al., 2025).
One other limitation of our study is that we rely exclusively on brain data collected from experiments conducted with English stimuli. As such, we do not explore whether our findings generalize across languages. This remains an open question and warrants further investigation. That said, evidence from cross-linguistic neuroscience research studying 45 languages from 12 language families (Malik-Moraleda et al., 2022) suggests the existence of a universal language network in the brain that is robust across languages and language families, both in topography and core functional properties.
Finally, a key question remains: Does LLM alignment evolution mirror human language acquisition? Comparing LLM representations to developmental data could reveal insights into learning trajectories and help differentiate formal from functional language learning. Expanding brain-scoring benchmarks and incorporating multimodal models will help address these questions, further bridging the gap between artificial and biological intelligence and deepening our understanding of how both systems process and represent language.
Ethical Statement
This research relies on previously published neuroimaging (fMRI, ECoG) and behavioral datasets, collected by the original research groups under their institutional ethical guidelines with informed consent and IRB/ethics approval. Our work involved only secondary analysis of de-identified data, with no new data collection or direct participant interaction, and we remain committed to using such data responsibly and respectfully.
Acknowledgments
We thank the members of the EPFL NeuroAI and NLP labs for their valuable feedback and insightful suggestions. We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant.